OpenStudy (lastdaywork):

Determinants - If f(x) is a polynomial of degree < 3 ; prove that

OpenStudy (lastdaywork):

 \[=\frac{ f(x) }{ (x-a)(x-b)(x-c) }\]

ganeshie8 (ganeshie8):

one observation is : for the denominator : a-b, b-c, and c-a are factors

ganeshie8 (ganeshie8):

not sure if thats useful, but i think we need to take cases when degree = 0, 1, 2 ?

OpenStudy (lastdaywork):

I found LHS \[=\frac{ f(a) }{ (x-a)(b-a)(c-a) }+\frac{ f(b) }{ (a-b)(x-b)(c-b) }+\frac{ f(c) }{ (a-c)(b-c)(x-c) }\] Donno what to do next.. :(

OpenStudy (lastdaywork):

Solution in the book says - "RHS will also be the same if we split it into partial fractions by the method of suppression by putting x=a,b,c successively." IDK what does that ^^ means..

ganeshie8 (ganeshie8):

thats just fancy way of saying "decompose RHS into partial fractions by using cover up method"

ganeshie8 (ganeshie8):

\(\large \frac{ f(x) }{ (x-a)(x-b)(x-c) } = \frac{A}{x-a} + \frac{B}{x-b} + \frac{C}{x-c}\)

ganeshie8 (ganeshie8):

since f(x) degree is less < 3, we can do that^

OpenStudy (anonymous):

\[ f(x) = k_2x^2+k_1x+k_0 \]plug an chug.

ganeshie8 (ganeshie8):

multiply thru by x-a, and cover up A : you wud get A = \(\frac{f(a)}{(a-b)(a-c)}\)

OpenStudy (lastdaywork):

Yea, I got it now..thanx ^_^

OpenStudy (lastdaywork):

BTW, what does "cover-up" means ?

ganeshie8 (ganeshie8):

watch this from 5:50 to 10:00

ganeshie8 (ganeshie8):

just an easy trick for doing partial fractions...

OpenStudy (lastdaywork):

Thanxx :)

ganeshie8 (ganeshie8):

np :)